# Modeling the Deterioration of Pavement Skid Resistance and Surface Texture After Preventive Maintenance

### v1

@@ Abstract
This study investigates the deterioration of skid resistance and surface macrotexture following preventive maintenance using micro-milling techniques. Field data were collected from 31 asphalt pavement sections located across four climatic zones in Texas, encompassing a variety of surface types, milling depths, operational speeds, and drum configurations. A standardized data collection protocol was followed, with measurements taken before milling, immediately after treatment, and at 3, 6, 12, and 18 months post-treatment. Skid number and Mean Profile Depth (MPD) were used to evaluate surface friction and texture characteristics. The dataset was reformatted into a time-series structure with 930 observations, incorporating contextual variables such as climatic zone, treatment parameters, and baseline surface condition. A comparative modeling framework was applied to predict the deterioration trends of both skid resistance and macrotexture over time. Eight regression models, including linear, tree-based, and ensemble methods, were evaluated alongside a sequence-to-one transformer model. Results show that the transformer model achieved the highest prediction accuracy for skid resistance ($R^{2}=0.981$), while Random Forest performing best for macrotexture prediction ($R^{2}=0.838$). The findings indicate that the degradation of surface characteristics after preventive maintenance is nonlinear and influenced by a combination of environmental and operational factors. This study demonstrates the effectiveness of data-driven modeling in supporting transportation agencies with pavement performance forecasting and maintenance planning.
Keywords: Pavement surface characteristics, skid resistance, surface texture, preventive maintenance, macrotexture, micro-milling, surface deterioration modeling, time-series pavement data, machine learning, asphalt pavement performance
@@ Introduction
Pavement surface characteristics such as skid resistance and macrotexture are crucial indicators for ensuring road safety and preventing accidents [1]. Skid resistance, the friction generated between tires and the pavement surface, is fundamental for vehicle control during critical maneuvers like turning and emergency braking [2]. A significant decrease in skid resistance can lead to a considerable increase in accident rates, and higher skid numbers are consistently associated with fewer crashes [3]. Macrotexture, defined as the surface texture component with wavelengths between 0.5 and 50 mm, plays a vital role in removing water film from the tire-pavement contact area and contributes to skid resistance, particularly at higher speeds and in wet conditions [4]. Both microtexture (roughness less than 0.5 mm) and macrotexture together determine the overall pavement friction. Pavement skid resistance undergoes a natural progression of change over time due to traffic and environmental influences: it may initially increase as the asphalt binder wears away, then decline as the aggregate becomes polished, and eventually reach a stable state [5]. Therefore, consistent monitoring and maintenance of these pavement surface characteristics, often quantified by the Skid Number and Mean Texture Depth, are essential to maintain adequate traction and mitigate crash risks, especially under wet weather conditions [6].
Preventive maintenance techniques, such as micro-surfacing and milling operations, are applied to significantly alter existing pavement surface properties [7]. For instance, micro-surfacing aims to improve anti-skid performance, wear resistance, and durability while also addressing noise reduction through optimized gradation and the inclusion of additives like rubber powder and fiber [8]. These operations are designed to modify surface characteristics like texture (e.g., macrotexture, ridge-to-valley depth) to enhance friction, drainage, and prepare the surface for subsequent overlays [9]. The post-treatment evolution of these altered properties becomes a key area of interest because factors such as the duration of exposure to traffic and weather after milling, the pavement temperature during milling, and the effectiveness of quality control directly influence the long-term performance, service life, and overall cost-effectiveness of the pavement and any new layers [10].
Given the significant role that preventive maintenance treatments play in modifying pavement surface properties, it becomes critically important to understand how these properties evolve after treatment. In particular, quantifying the effectiveness and deterioration rate of surface characteristics such as skid resistance and macrotexture following preventive treatments is essential for ensuring long-term safety and performance. The post-treatment behavior of these indicators can vary widely depending on factors such as surface type, climatic conditions, and operational parameters used during the treatment process. Without a clear understanding of these dynamics, it is challenging to optimize treatment timing, select appropriate technologies, or accurately forecast service life and maintenance needs.
Given the critical role of pavement surface characteristics, particularly skid resistance and macrotexture, in ensuring road safety, there is a pressing need to understand how these properties evolve following preventive maintenance treatments. While micro-milling has emerged as a widely adopted technique to restore surface friction and texture, the long-term effectiveness of such treatments remains insufficiently quantified. Existing models often fail to capture the temporal dynamics of surface deterioration, treating observations as independent and overlooking the influence of prior conditions. This study addresses the following research question: How do skid resistance and surface macrotexture deteriorate over time after micro-milling treatments, and to what extent can sequence-based machine learning models improve the prediction of this degradation compared to traditional approaches? By framing the deterioration process as a temporal learning problem, this research aims to enhance the predictive accuracy of pavement performance models and support more informed maintenance planning.
@@
A strong correlation exists between pavement surface features and friction values, with higher skid resistance directly linked to improved traction and reduced crash frequencies, particularly under wet conditions [3]. Studies have shown that a 10-unit increase in Skid Number (SN) can lead to a 7% to 8% reduction in dry-weather crashes and a more significant 14% to 21% decrease in wet-weather crashes on highways [17]. The impact of increased skid resistance is especially pronounced on interstate highways and curved segments during wet conditions, where friction demand is higher [18]. Furthermore, vehicle dynamics simulations indicate that factors such as driving speed, vehicle type, and polishing time significantly influence critical safety metrics like braking distance and lateral excursion distance [19].
@@
The effectiveness of micro-milling on surface texture and skid resistance is significantly influenced by operational factors. Sections milled with fine drums, typical of micro-milling, demonstrate enhanced skid resistance and macrotexture [7]. Higher forward milling speeds also correlate positively with improved skid resistance and macrotexture, with recommended speeds ranging from 70 to 80 feet per minute [9]. While optimal cutting depths for texturing are typically between 0.25 and 0.5 inches (6 to 13 mm), the milling depth itself does not notably impact the long-term deterioration of skid resistance or macrotexture [21]. To ensure proper bonding with subsequent overlays, a clean milled surface, often achieved with power brooms and brief traffic exposure to clear dust, is essential [7].
Despite its benefits, micro-milling presents challenges, particularly in achieving uniform texture and depth, which can be affected by factors like original pavement distress, machine performance, and drum rotation rates [22]. Inconsistent milling can compromise bonding strength, leading to premature distresses such as peeling or shoving in the overlay [23]. Micro-milling has demonstrated service lives ranging from about 12 months on seal coats to 10-12 years on hot-mix asphalt (HMA) sections [24].
@@
Pavement skid resistance is not static; it continuously deteriorates over time due to various influencing factors, primarily traffic loads and environmental conditions [25]. This deterioration typically progresses through stages: an initial rapid increase (due to binder removal), followed by a slow decrease as aggregate micro-texture worsens, and finally a stabilization phase [26]. Vehicular traffic, through its polishing effect, is a major contributor to this decline, especially in the wheel path areas [27]. The mode of vehicle movement dictates the directionality of polishing; for instance, acceleration-dominated sections experience higher friction against traffic flow, while deceleration-dominated sections show lower friction in the direction of traffic flow [28]. Environmental factors such as the presence of water, snow, ice, and other contaminants like rubber deposits significantly reduce skid resistance by forming a lubricating film or covering the texture [29]. Water film thickness, in particular, negatively correlates with skid resistance and can lead to hydroplaning at higher speeds. Temperature (both tire and ambient/pavement) can also influence skid resistance, with increases often leading to a reduction [30].
Beyond traffic and environmental effects, specific material and design choices significantly impact skid resistance and its long-term performance [31]. The mineralogy of aggregates greatly influences their susceptibility to wear and polishing; for example, poly-mineral aggregates tend to exhibit higher polishing resistance and friction than mono-mineral ones like limestone [32]. Aggregate dimension also plays a role, with finer aggregates typically providing higher initial friction but experiencing a more pronounced reduction with polishing [33]. For asphalt mixtures, the long-term skid resistance is primarily determined by coarse aggregates, while fine aggregates mainly affect the initial friction [34]. The type and dosage of binder can influence the initial stripping phase, affecting both friction values and the time required for complete binder removal [35]. Pavement macro-texture and air voids are crucial, as they influence the actual tire-pavement contact area and drainage capacity [36]. The use of recycled materials like reclaimed asphalt pavement (RAP) or crumb rubber, and warm mix asphalt (WMA) additives, may have a detrimental effect on skid resistance and can delay binder removal [37]. Grooving pavements is an effective measure to enhance drainage and skid resistance, and reduce hydroplaning risk, but groove deterioration due to wear or rubber deposits can compromise their effectiveness [38]. Pavement markings also affect skid resistance, with marking materials and glass beads potentially reducing friction compared to the surrounding pavement, though anti-skid additives can mitigate this [39].
@@
The modeling of pavement skid resistance and texture deterioration is crucial for road safety and maintenance, addressing the complex interactions between vehicle tires and road surfaces [40]. Given the challenges of direct and continuous field measurements, various analytical, statistical, and machine learning models have been developed to predict and understand these time-dependent properties [41, 42, 43].
@@
Statistical and analytical models have traditionally been used to describe the decay of pavement conditions [40, 44]. Common approaches include empirical relationships and regression models that link pavement characteristics, such as texture parameters (Mean Texture Depth (MTD) and Mean Profile Depth (MPD)), to skid resistance or its rate of decline [hu2022evaluate]. Models like the Penn State model, PIARC International Friction Index (IFI), and the Rose and Gallaway model are examples that estimate skid resistance as a function of speed and texture, though they often have limitations regarding water film thickness or application scope [45]. The observed decay law of skid resistance often shows an initial rapid decline, followed by a slower decrease, eventually stabilizing, which is captured by models such as the asymptotic, exponential, and logarithmic models [46]. However, these statistical methods may struggle with non-linear relationships and the multitude of influencing factors, leading to limited accuracy and applicability [47].
@@
To overcome the limitations of traditional statistical models, machine learning (ML) has emerged as a powerful tool for predicting skid resistance and texture, capable of handling complex, non-linear relationships and multi-dimensional data [48]. Widely applied ML algorithms include Artificial Neural Networks (ANN), Support Vector Machines (SVM), Random Forest (RF), and Decision Trees (DT) [48]. These models leverage various input parameters, such as vehicle speed, water depth, tire tread depth, pavement macro-texture, aggregate properties, and even environmental factors, to predict friction coefficients or classify skid resistance levels (e.g., low, medium, high) [49]. Data from 3D laser scanning and other non-contact measurement methods provide high-resolution texture information as input for these ML models, improving prediction accuracy [12].
@@
More advanced AI and ensemble learning models are continually being developed to enhance prediction accuracy and interpretability [50, 51]. Back Propagation (BP) neural networks, often optimized by genetic algorithms (GA-BP), have shown strong self-learning abilities and can establish highly non-linear relationships for skid resistance prediction, even for long-term analysis or specific pavement types like steel bridge deck pavement [52]. Light Gradient Boosting Machine (LightGBM) and its optimized variant, Bayesian-LightGBM, are noted for their stability, generalization ability, and efficiency in handling multi-factor regression problems, significantly improving prediction accuracy compared to other models [hu2022evaluate]. Deep learning approaches, such as Convolutional Neural Networks (CNN) and deep residual networks (ResNets), are also being explored, particularly with 3D texture data, to achieve higher accuracy in correlating texture with friction [12, 53]. Integrated frameworks combining techniques like Fast Fourier Transform (FFT) with Extreme Gradient Boosting (XGBoost) or wavelet transforms with fractal theory further refine texture characterization and improve skid resistance prediction, offering insights into the impact of multi-scale texture features [54].
While these models offer significant improvements, challenges remain. The accuracy of ML/AI models is highly dependent on the quality and quantity of data. These models with missing values, errors, or unrepresentative samples often lead to deviations in predicted results. Many models also do not fully account for the complex coupled effects of multiple factors (e.g., traffic, climate, material interactions) or lack real-time update capabilities. Future research directions emphasize expanding datasets, developing multivariate and high-precision combined prediction models, and incorporating dynamic changes in road conditions, temperature, humidity, and the interaction mechanisms of micro- and macro-textures to create more robust, reliable, and practically effective models for long-term skid resistance assessment.
@@
Predicting the temporal evolution of pavement skid resistance is a classic sequence–to–one regression task. For a given test section, we observe a finite sequence of feature vectors $\mathbf{x}_{0},\mathbf{x}_{1},\dots,\mathbf{x}_{T}$ recorded at discrete time points $t_{0},t_{1},\dots,t_{T}$, where each $\mathbf{x}_{t}\in\mathbb{R}^{d_{x}}$ represents the concatenation of all feature values observed at time $t$. Given a sliding window of length $L$, the Transformer receives the ordered sequence
@@
The Transformer is a neural architecture designed for sequence modeling. The Transformer architecture processes all inputs within a common latent space of fixed dimensionality $d_{\text{model}}$. This design allows the model to uniformly apply self-attention and feed-forward layers regardless of the input’s original structure or scale. To accommodate this, each input vector $\mathbf{x}_{t}\in\mathbb{R}^{d_{x}}$ is first projected into the latent space through a learned linear transformation:
@@
To retain the temporal order of the sequence, a positional encoding vector $\mathbf{p}_{t}\in\mathbb{R}^{d_{\text{model}}}$ is added to each embedding. This positional signal distinguishes tokens from different time steps, even if their feature content is similar. A widely used deterministic choice is the sinusoidal encoding:
@@
Let $\mathbf{H}^{(l)}\in\mathbb{R}^{L\times d_{\text{model}}}$ be the input to layer $l$. The self-attention mechanism computes query, key, and value matrices:
@@
Using $H$ independent attention heads allows the model to capture diverse interaction patterns. Their outputs are concatenated and linearly projected back:
@@
A feed-forward network is then applied to each time step independently:
@@
and the final output of the encoder layer is:
@@
Because we require a single prediction, we extract the final time-step embedding $\mathbf{h}_{t}^{(N)}$ and feed it to a small multi-layer perceptron:
@@
The network parameters $\Theta$ are trained to minimise the mean-squared error
@@
This section compares the performance of eight traditional machine–learning baselines with a Sequence-to-One Transformer Regression Model. The Transformer takes a sliding window of length $L$ (here $L=4$) and predicts the skid number or macrotexture value at the next inspection month. For fair comparison, all models were evaluated on the same train–test split (80 % / 20 %) and assessed with the coefficient of determination ($R^{2}$), root-mean-squared error (RMSE), and mean-absolute error (MAE).
Table 2 shows that the Transformer achieves the highest accuracy for skid-number forecasting ($R^{2}=0.981$), edging out XGBoost ($R^{2}=0.979$) and outperforming all other baselines by a clear margin. Tree-based models (Random Forest, Decision Tree) rank next, whereas linear regressions and the MLP network trail far behind, indicating that non-linear interactions and temporal dependencies are critical for capturing friction degradation.
@@
Across both targets, ensemble trees (Random Forest, XGBoost) and the Transformer dominate, confirming that non-linear effects, feature interactions, and temporal patterns are essential for accurate forecasting. The Transformer’s slight advantage on skid number demonstrates the value of explicitly modeling sequential dependencies, whereas its comparable performance on macrotexture suggests diminishing returns when historical texture inherently exhibits weaker temporal autocorrelation. These results highlight the practicality of the Sequence-to-One Transformer for proactive pavement-management applications, while also showing that well-tuned ensemble methods remain competitive baselines when computational simplicity is desired.

### v2

# Time Series Transformer-Based Modeling of Pavement Skid and Texture Deterioration

@@ Abstract
This study investigates the deterioration of skid resistance and surface macrotexture following preventive maintenance using micro-milling techniques. Field data were collected from 31 asphalt pavement sections located across four climatic zones in Texas. The data encompasses a variety of surface types, milling depths, operational speeds, and drum configurations. A standardized data collection protocol was followed, with measurements taken before milling, immediately after treatment, and at 3, 6, 12, and 18 months post-treatment. Skid number and Mean Profile Depth (MPD) were used to evaluate surface friction and texture characteristics. The dataset was reformatted into a time-series structure with 930 observations, including contextual variables such as climatic zone, treatment parameters, and baseline surface condition. A comparative modeling framework was applied to predict the deterioration trends of both skid resistance and macrotexture over time. Eight regression models, including linear, tree-based, and ensemble methods, were evaluated alongside a time series transformer model. Results show that the transformer model achieved the highest prediction accuracy for skid resistance ($R^{2}=0.981$), while Random Forest performing best for macrotexture prediction ($R^{2}=0.838$). The findings indicate that the degradation of surface characteristics after preventive maintenance is nonlinear and influenced by a combination of environmental and operational factors. This study demonstrates the effectiveness of data-driven modeling in supporting transportation agencies with pavement performance forecasting and maintenance planning.
Keywords: Pavement deterioration, skid resistance, pavement performance modeling, transformer, time series, macrotexture, micro-milling, machine learning, deep learning, preventive maintenance
@@ Introduction
Pavement surface characteristics such as skid resistance and macrotexture are crucial indicators for ensuring road safety and preventing accidents [1]. Skid resistance, the friction generated between tires and the pavement surface, is fundamental for vehicle control during critical maneuvers like turning and emergency braking [2]. A significant decrease in skid resistance can lead to a considerable increase in accident rates, and higher skid numbers are consistently associated with fewer crashes [3]. Macrotexture, defined as the surface texture component with wavelengths between 0.5 and 50 mm, plays a vital role in removing water film from the tire-pavement contact area and contributes to skid resistance, particularly at higher speeds and in wet conditions [4]. Both microtexture (roughness less than 0.5 mm) and macrotexture together determine the overall pavement friction. Pavement skid resistance undergoes a natural progression of change over time due to traffic and environmental influences. It may initially increase as the asphalt binder wears away, then decline as the aggregate becomes polished, and eventually reach a stable state [5]. Therefore, consistent monitoring and maintenance of these pavement surface characteristics are essential to maintain adequate traction and mitigate crash risks, especially under wet weather conditions [6].
Preventive maintenance techniques, such as micro-surfacing and milling operations, are applied to significantly alter existing pavement surface properties [7]. For instance, micro-surfacing aims to improve anti-skid performance, wear resistance, and durability. It also addresses noise reduction through optimized gradation and the inclusion of additives like rubber powder and fiber [8]. However, the effectiveness of such treatments remains insufficiently quantified. This study addresses the following research question: How do skid resistance and surface macrotexture deteriorate over time after micro-milling treatments; and to what extent can sequence-based machine learning models improve the prediction of this degradation compared to traditional approaches?
@@
A strong correlation exists between pavement surface friction values and road safety. Higher skid resistance usually linked to reduced crash frequencies, particularly under wet conditions [3]. Studies have shown that a 10-unit increase in Skid Number (SN) can lead to a 7% to 8% reduction in dry-weather crashes and a more significant 14% to 21% decrease in wet-weather crashes on highways [15]. The impact of increased skid resistance is especially pronounced on interstate highways and curved segments during wet conditions, where friction demand is higher [16].
@@
The effectiveness of micro-milling on surface texture and skid resistance is significantly influenced by operational factors. Sections milled with fine drums, typical of micro-milling, demonstrate enhanced skid resistance and macrotexture [7]. Higher forward milling speeds also correlate positively with improved skid resistance and macrotexture. The recommended speeds ranges from 70 to 80 feet per minute [18]. Optimal cutting depths for texturing are typically between 0.25 and 0.5 inches (6 to 13 mm). However, the milling depth itself does not notably impact the long-term deterioration of skid resistance or macrotexture [20].
Micro-milling faces challenges in achieving consistent texture and depth. Factors such as existing pavement distress, machine performance, and drum rotation rates can influence these outcomes [21]. Inconsistent milling can compromise bonding strength, leading to premature distress such as peeling or shoving in the overlay [22]. Micro-milling has demonstrated service lives ranging from about 12 months on seal coats to 10-12 years on hot-mix asphalt (HMA) sections [23].
@@
Pavement skid resistance is not static. It continuously deteriorates over time due to various influencing factors, primarily traffic loads and environmental conditions [24]. The deterioration process usually occurs in stages: first, a rapid increase due to binder removal, then a slow decrease as aggregate micro-texture worsens, and finally, a stabilization phase [25]. Vehicular traffic, through its polishing effect, is a major contributor to this decline, especially in the wheel path areas [26]. Vehicle movement determines the direction of polishing. Acceleration-dominated areas face higher friction against traffic flow, while deceleration-dominated areas experience lower friction in the direction of traffic flow [27]. Environmental factors, including water, snow, ice, and contaminants like rubber deposits, significantly reduce skid resistance. They do this by forming a lubricating film or covering the surface texture [28]. Water film thickness, in particular, negatively correlates with skid resistance and can lead to hydroplaning at higher speeds. Temperature (both tire and ambient/pavement) can also influence skid resistance, with increases often leading to a reduction [29].
Beyond traffic and environmental effects, specific material and design choices, such as the use of composite aggregates, significantly impact skid resistance and its long-term performance [30, 31]. The mineralogy of aggregates greatly influences their susceptibility to wear and polishing. For example, poly-mineral aggregates tend to exhibit higher polishing resistance and friction than mono-mineral ones like limestone [32]. Aggregate dimension also plays a role, with finer aggregates typically providing higher initial friction but experiencing a more pronounced reduction with polishing [33]. For asphalt mixtures, the long-term skid resistance is primarily determined by coarse aggregates, while fine aggregates mainly affect the initial friction [34].
The type and dosage of binder can influence the initial stripping phase, which affects both friction values and the time required for complete binder removal [35]. Pavement macro-texture and air voids are crucial, as they influence the actual tire-pavement contact area and drainage capacity [36]. The use of recycled materials like reclaimed asphalt pavement (RAP) or crumb rubber, and warm mix asphalt (WMA) additives, may have a detrimental effect on skid resistance and can delay binder removal [37]. Grooving pavements improves drainage and skid resistance, and reduces hydroplaning risk. However, groove deterioration from wear or rubber deposits can lessen their effectiveness [38]. Pavement markings also influence skid resistance. The materials used and glass beads can decrease friction compared to the surrounding pavement. However, anti-skid additives can help to reduce this effect. [39].
@@
Modeling pavement skid resistance and texture deterioration is vital for road safety and maintenance. It helps understand the complex interactions between vehicle tires and road surfaces [40]. Various analytical, statistical, and machine learning models have been developed to predict and understand these time-dependent properties [41, 42, 43].
@@
Statistical and analytical models have traditionally been used to describe the decay of pavement conditions [44, 45, 46]. Common approaches include empirical relationships and regression models. These link pavement characteristics to skid resistance or its decline rate. Characteristics used include texture parameters like Mean Texture Depth (MTD) and Mean Profile Depth (MPD) [47]. Models like the Penn State model, PIARC International Friction Index (IFI), and the Rose and Gallaway model are examples that estimate skid resistance as a function of speed and texture, though they often have limitations regarding water film thickness or application scope [48]. Skid resistance often declines rapidly at first, then slows, eventually stabilizing. Models like asymptotic, exponential, and logarithmic ones capture this behavior [49]. However, these statistical methods may struggle with non-linear relationships and the multitude of influencing factors, leading to limited accuracy and applicability [50].
@@
Machine learning (ML) has emerged as a powerful tool for predicting pavement conditions. It overcomes the limitations of traditional statistical models by handling complex, non-linear relationships and multi-dimensional data [51, 52, 53, 54, 55]. Widely applied ML algorithms include Artificial Neural Networks (ANN), Support Vector Machines (SVM), Random Forest (RF), and Decision Trees (DT) [51]. These models leverage various input parameters, such as vehicle speed, water depth, tire tread depth, pavement macro-texture, aggregate properties, and even environmental factors, to predict friction coefficients or classify skid resistance levels (e.g., low, medium, high) [56]. Data from 3D laser scanning and other non-contact measurement methods provide high-resolution texture information as input for these ML models, improving prediction accuracy [10].
@@
More advanced AI and ensemble learning models are continually being developed to enhance prediction accuracy and interpretability [57, 53]. Neural networks optimized by genetic algorithms demonstrate strong self-learning capabilities. They can establish highly non-linear relationships for skid resistance prediction. This applies even to long-term analysis and specific pavement types, such as steel bridge deck pavement [58]. Light Gradient Boosting Machine (LightGBM) and its optimized variant, Bayesian-LightGBM, are known for their stability, generalization, and efficiency in multi-factor regression. They significantly improve prediction accuracy over other models [47].
Deep learning approaches, such as Convolutional Neural Networks (CNN) and deep residual networks (ResNets), are also being explored, particularly with 3D texture data, to achieve higher accuracy in correlating texture with friction [10, 59]. Integrated frameworks that combine techniques such as Fast Fourier Transform (FFT) with Extreme Gradient Boosting (XGBoost) or wavelet transforms with fractal theory enhance texture characterization and improve skid resistance prediction. These frameworks provide insights into the impact of multi-scale texture features [60].
@@
Problem Formulation
Predicting the temporal evolution of pavement skid resistance is a classic sequence–to–one time series regression task. In this task, the model uses a sequence of past pavement conditions to predict what the road will be like at the next inspection. For example, if the past five years of data are available, the model learns patterns from those five years and uses them to forecast the skid resistance and texture for the sixth year. For a given test section, a finite sequence of feature vectors $\mathbf{x}_{0},\mathbf{x}_{1},\dots,\mathbf{x}_{T}$ is observed at discrete time points $t_{0},t_{1},\dots,t_{T}$, where each $\mathbf{x}_{t}\in\mathbb{R}^{d_{x}}$ represents the concatenation of all feature values observed at time $t$. Given a sliding window of length $L$, the Transformer receives the ordered sequence
@@
To capture both short- and long-term temporal dependencies in this sequence-to-one prediction task, the Transformer model [61] was adopted. Its self-attention mechanism makes it well-suited for time series forecasting, especially when handling long-range dependencies and multivariate inputs.
Time Series Transformer
The Transformer architecture, originally developed for natural language processing, is well suited for time series forecasting tasks with complex temporal dependencies. Unlike traditional models such as ARIMA or recurrent neural networks (RNNs), which process sequences sequentially, the transformer utilizes a self-attention mechanism that considers all time steps simultaneously. This enables the model to more effectively capture long-range dependencies and interactions among variables across time [62].
Time Series transformers adapt the original transformer architecture to forecasting problems by replacing word embeddings with time series feature vectors and modifying the positional encoding to reflect temporal structure. These models are capable of handling multivariate inputs, irregular time steps, and long input sequences. Time series transformers have been widely applied in multivariate forecasting tasks [63, 64, 65]. The following section introduces the key concepts of the Transformer architecture.
@@
The Transformer is a neural architecture designed for sequence modeling. The Transformer architecture processes all inputs within a common latent space of fixed dimensionality $d_{\text{model}}$. This design allows the model to uniformly apply self-attention and feed-forward layers regardless of the input’s original structure or scale. To accommodate this, each input vector $\mathbf{x}_{t}\in\mathbb{R}^{d_{x}}$ is first projected into the latent space through a learned linear transformation [61]:
@@
To retain the temporal order of the sequence, a positional encoding vector $\mathbf{p}_{t}\in\mathbb{R}^{d_{\text{model}}}$ is added to each embedding. This positional signal distinguishes tokens from different time steps, even if their feature content is similar. A widely used deterministic choice is the sinusoidal encoding [61]:
@@
Let $\mathbf{H}^{(l)}\in\mathbb{R}^{L\times d_{\text{model}}}$ be the input to layer $l$. Following the scaled‑dot attention of Vaswani et al. [61], the self-attention mechanism computes query, key, and value matrices:
@@
Using $H$ independent attention heads allows the model to capture diverse interaction patterns. Their outputs are concatenated and linearly projected back [61]:
@@
Following Vaswani et al. [61], a feed-forward network is then applied to each time step independently:
@@
Using residual connections [66] and layer normalization [67], the final output of the encoder layer is:
@@
Following common practice in Transformer-based regression models (e.g., Vaswani et al. [61]), the final time-step embedding $\mathbf{h}_{t}^{(N)}$ is extracted and passed through a small multi-layer perceptron.
@@
The network parameters $\Theta$ are trained to minimise the mean-squared error, commonly used in regression tasks [68].
@@
This section compares the performance of eight traditional machine–learning baselines with a time series transformer regression Model. The Transformer takes a sliding window of length $L$ (here $L=4$) and predicts the skid number or macrotexture value at the next inspection month. For fair comparison, all models were evaluated on the same train–test split (80 % / 20 %) and assessed with the coefficient of determination ($R^{2}$), root-mean-squared error (RMSE), and mean-absolute error (MAE).
The time series transformer model is implemented in PyTorch and consists of two stacked encoder layers with four self-attention heads and a hidden dimension of 64. Input features at each time step are projected into a 64-dimensional latent space, followed by the addition of learnable positional encodings. The output from the last time step is passed through a two-layer feed-forward network with a hidden size of 32 and ReLU activation to produce a scalar prediction. The model is trained using the Adam optimizer with a learning rate of $10^{-3}$ and a batch size of 16, using mean squared error as the loss function. Table 2 shows that the Transformer achieves the highest accuracy for skid-number forecasting ($R^{2}=0.981$), edging out XGBoost ($R^{2}=0.979$) and outperforming all other baselines by a clear margin. Tree-based models (Random Forest, Decision Tree) rank next, whereas linear regressions and the MLP network trail far behind, indicating that non-linear interactions and temporal dependencies are critical for capturing friction degradation.
@@
Across both targets, ensemble trees (Random Forest, XGBoost) and the Transformer dominate, confirming that non-linear effects, feature interactions, and temporal patterns are essential for accurate forecasting. The Transformer’s slight advantage on skid number demonstrates the value of explicitly modeling sequential dependencies, whereas its comparable performance on macrotexture suggests diminishing returns when historical texture inherently exhibits weaker temporal autocorrelation. These results highlight the practicality of the time series transformer for proactive pavement-management applications, while also showing that well-tuned ensemble methods remain competitive baselines when computational simplicity is desired.